### v1

# Towards Quantum Advantage in Chemistry

Abstract
Molecular simulations are widely regarded as leading candidates to demonstrate quantum advantage–defined as the point at which quantum methods surpass classical approaches in either accuracy or scale. Yet the qubit counts and error rates required to realize such an advantage remain uncertain; resource estimates for ground-state electronic structure span orders of magnitude, and no quantum-native method has been validated at a commercially relevant scale. Here we address this uncertainty by executing the iterative qubit coupled-cluster (iQCC) algorithm, designed for fault-tolerant quantum hardware, at unprecedented scale using a quantum solver on classical processors, enabling simulations of transition organo-metallic complexes requiring hundreds of logical qubits and millions of entangling gates. Using this approach, we compute the lowest triplet excited state (T1) energies of Ir(III) and Pt(II) phosphorescent organometallic compounds and show that iQCC achieves the lowest mean absolute error (0.05 eV) and highest R2 (0.94) relative to experiment, outperforming leading classical methods. We find these systems remain classically tractable up to 200 logical qubits, establishing the threshold at which quantum advantage in computational chemistry may emerge and clarifying resource requirements for future quantum computers.
1 Introduction
Predicting molecular properties with high accuracy remains a central challenge in chemistry, materials, and drug discovery [1, 2, 3]. Conventional simulation approaches such as density-functional theory (DFT) are widely used but limited in accuracy [4, 1, 5], while the gold standard coupled-cluster methods achieve chemical precision only at prohibitive computational cost [6]. In principle, quantum computers promise to overcome these barriers by natively representing electronic wavefunctions [7, 8, 9]. Recent advances have extended the range and now allow classical computers to emulate the variational quantum eigensolver (VQE) algorithm in the 80–92 qubit range by using, for example, a sparse representation of the wavefunction [10], tensor networks on super-computer clusters [11, 12], or combined hybrid quantum solvers [13], but current devices lack the fidelity and scale required for commercially relevant systems [14, 15, 16].
Here we present advances in a Quantum Solver that executes the iterative qubit coupled cluster (iQCC) quantum algorithm [17, 18, 19] at an unprecedented scale, enabling direct benchmarking of quantum-native approaches against both experiment and state-of-the-art classical methods. By enhancing the efficiency of the iQCC Quantum Solver, we simulate problems equivalent to 200 logical qubits and two-qubit gates, far beyond previous emulations and current quantum hardware [14, 11, 20, 13, 16].
As a benchmark, we compute excited state energies of phosphorescent iridium (III) and platinum (II) complexes used in organic light-emitting diodes (OLEDs) [21, 22, 23, 24]. These complexes combine industrial importance with experimentally accessible properties, making them ideal for assessing accuracy. Our results demonstrate systematic improvements over conventional methods and establish a framework for evaluating quantum advantage in chemistry, materials, and drug discovery.
2 Results
2.1 Summary of iQCC
The iQCC method is a VQE type quantum algorithm which uses a unitary ansatz and represents a trial wave function as
where is an initial state of a quantum register, typically a one-determinant representation of the ground-state wave function written in a population basis, and is a product of factors , where are products of the Pauli matrices. This qubit coupled cluster (QCC) form, unlike other commonly used forms, such as the unitary coupled cluster singles and doubles (UCCSD), derives its generators (“entanglers”) directly from a qubit Hamiltonian to guarantee lowering of the total ground-state energy. This avoids the barren plateau problem that other implementations struggle with [17, 18] and does not require amplitudes to be solved using coupled-cluster singles and doubles (CCSD) then inputted as an initial guess [25]. Amplitude optimization is performed using the exact or an approximate ansatz [26] by computing the energy of the Hamiltonian:
where of which are real coefficients derived from one- and two-electron integrals and are Pauli words. In the iQCC method, the original Hamiltonian is unitary-transformed (“dressed”) using the optimized ansatz into the subsequent iQCC Hamiltonian (Fig. 1a). At each iQCC iteration, a subset of the direct interaction set (DIS) (the set of all possible entanglers with non-zero gradients) is included in the QCC ansatz, selected to include the set of all Pauli terms which can be computed efficiently classically. To approximately account for the energy contributions of entanglers from the DIS that are not included into the QCC ansatz, we apply the Epstein-Nesbet perturbation theory (PT) correction [19], to obtain the iQCC +PT energy estimate.
2.2 Scalability Analysis and Benchmarking
To enable large‑scale classical simulations of iQCC, we implemented the algorithm in C++ with OpenMPI, parallelizing the most expensive components—operator dressing and amplitude optimization (Fig. 1a). The approach uses a bit-partitioning scheme where each Pauli word is represented in symplectic binary form, and selected bits serve as keys that determine CPU assignment. This binary encoding not only yields linear memory scaling, with respect to the number of qubits and Pauli words in the Hamiltonian, but also provides an explicit, computation-free mapping of every Pauli word to its designated processor [27] (full details given in Methods Section 4.3). During the optimization stage, the method supports complete optimization of the exact QCC ansatz [18, 28] as well as polynomial approximate forms [26]. Some key advantages for benchmarking chemistry simulations using the iQCC quantum solver provides are: 1) full qubit-to-qubit connectivity, 2) the energy can be directly computed without noise giving consistent and reliable results (see Methods 4.1 for further details), 3) the QCC anstaz has a direct circuit equivalent that can be run on other simulators or quantum hardware making it quantum-native [28, 29, 10] (details given in Supplementary Information Section SI.4 Quantum Circuit Equivalents), and 4) the solver has direct analytical access to the gradient for each parameter in the anstaz. This allows the quantum solver to use derivative based optimization, greatly improving it’s efficiency over quantum hardware implementations [30].
The iQCC algorithm has several advantages over other quantum simulation methods such as state-vector and tensor networks. State-vector methods require storing the full amplitude set in , which causes both memory and computational cost to scale as with respect to the number of qubits [31, 32]. Tensor network approaches reduce this burden by factorizing into low-rank tensors at cost , where is the bond dimension and depends on network topology (typically 3-4), but they remain efficient only when entanglement follows an area law; under volume law entanglement grows rapidly and the cost can approach or worse [33, 34, 35, 36, 37]. Instead of explicitly representing the wavefunction, the iQCC algorithm works in the operator space and gives direct access to Hamiltonian coefficients that quantify the contribution of each Pauli term to the energy. This distinction carries over to energy evaluation: hardware-based or variational methods estimate by statistically sampling from repeated measurements, whose sampling requirements grow rapidly with system size. iQCC on the other hand, has deterministic access to , and their dressed counterparts on a classical computer, permitting exact energy evaluation at each iteration, stable convergence, and precise control over accuracy and computational effort. This efficient parallel iQCC implementation allows molecular systems as large as 200 qubits to be simulated with their variational estimate reliably computed below the CISD energy using less than 800 GB of RAM (Fig. 1c, 1d and Table 1).
The strong scaling study of a 36 qubit Hamiltonian of water (complete active space (CAS) of 8 electrons and 18 orbitals – CAS(8,18), Å, basis 6-31G(d) – shows an excellent speedup from 1 to 32 processes, tapering off beyond 64 processes (Fig. 1b). The ansatz optimization steps, which include the construction of the Hessian for amplitude () optimization (Fig. 1b blue), benefit less compared to the dressing steps (Fig. 1b red). The final energy for water produced by simulating a 2000 entangler QCC ansatz was -76.19640996 Eh, which is within sub-milliHartree agreement with the literature value of -76.196411 Eh [26].
In addition to efficient storage and speedup, we further demonstrate that the total wall clock time of the iQCC quantum solver is generally linear with respect to qubit count when considering electronic Hamiltonians between systems from 80 to 200 qubits (Fig. 1c, Table SI.2-2). To illustrate this, we also normalize the scaling reported in Fig. 1c with respect to number of final entanglers to be optimized at each qubit count; this demonstrates that the main contributor to the total run time is the non-linear scaling of entangler optimization due to the increase in the required number of parameters to surpass the CISD energy. Considering that other quantum simulations scale as [33, 34, 35, 36, 37] or even [31, 32], this performance is outstanding within the scope of the quantum computing simulations currently available.
| Qubit Count | Number of Hamiltonian Terms [109] | Maximum Hamiltonian Storage [GB] | Number of Entanglers | 2-qubit CNOT Gates [106] | Variational iQCC Energy [Eh] | CISD Energy [Eh] |
| 80 | 2.48 | 398.2 | 300,000 | 2.859 | -1938.06200 | -1938.05499 |
| 100 | 2.87 | 423.2 | 400,000 | 3.398 | -1938.08067 | -1938.07434 |
| 140 | 2.92 | 537.1 | 600,000 | 4.300 | -1938.11184 | -1938.10940 |
| 160 | 2.99 | 540.0 | 800,000 | 5.597 | -1938.12528 | -1938.12328 |
| 200 | 3.75 | 755.8 | 1,500,000 | 10.210 | -1938.16627 | -1938.16010 |
Finally, we consider the performance of iQCC in the context VQE. The set of entanglers () optimized by the iQCC quantum solver have a direct mapping to quantum circuits (Supplementary Information Section SI.4 Quantum Circuit Equivalents). The iQCC quantum solver is capable of simulating circuits that have 200 logical qubits (CAS(100,100)) while optimising 1.5 parameters, which corresponds to an abstract circuit that contains 2-qubit CNOT gates (Table 1). To date, no universal quantum computer or simulator has executed this number of 2-qubit CNOT gates or qubits for an electronic structure calculation.
2.3 Quantum Utility Study: Estimating the peak PL emission of OLED materials
To assess the accuracy of the iQCC quantum solver in realistic systems, we selected a benchmark set of fourteen Ir (III) and Pt (II) phosphorescent complexes relevant to OLED s as shown in Fig. 2 [23, 24, 38, 39, 40]. Asymmetric Ir (III) complexes were chosen to complement prior studies on symmetric emitters, and the Pt (II) emitters were selected based on well-established scaffolds [23, 24, 40]. These materials span a wide range of phosphorescent emission energies from the deep blue of 440 nm to orange at 630 nm.
Experimental photoluminescence (PL) spectra for twelve of the materials were measured at 77 K (details in Methods Section 4.6) using toluene, THF, and dichloromethane which have relatively lower dielectric constants. The PL spectra for materials Q12 and Q14 was obtained from literature which were reported measured at 77 K in 2-MeTHF [41, 24]. The low dielectric solvent environment suppresses chromatic shifts, while cryogenic temperatures reduce vibrational broadening [28]. We omit spin orbit coupling since past work on similar compounds found it was not significant [42]. Thus, discrepancies between theoretical predictions and experiment are unlikely to arise from solvation or temperature and most likely arise from basis set error, allowing a more direct assessment of the intrinsic accuracy of the quantum and classical methods compared here.
Within the OLED industry, it is quite common to simulate the T S0 gap using either time-dependent density-functional theory (TD-DFT) or DFT, where the B3LYP and CAM-B3LYP functionals have been generally used [22, 41, 24, 28]. Past studies on multiple DFT functionals demonstrated that long range corrected hybrid functionals with HF exchange correlation, performed the best compared to hybrid functionals [28]. Since we seek to investigate whether quantum advantage can be gained by computing energies of molecular systems, it is important to compare the results against high accuracy methods widely used as benchmarks such as CCSD and completely renormalized coupled-cluster with singles, doubles and perturbative triples correction (CR-CC(2,3)) [43, 44, 45, 46, 47]. After performing the CR-CC(2,3) simulations and analyzing the diagnostic amplitude (details in Supplementary Information Section SI.5-1)[48], it was determined the materials were not strongly correlated, which means that methods such as density-matrix renormalization group (DMRG) would not provide a significant advantage over CR-CC(2,3). These ab initio methods have polynomial scaling and have been provided as examples as to why quantum advantage would be difficult to obtain for ground state simulation of weakly to moderately correlated chemical systems [15].
A computational chemistry tool needs to be considered precise in order to predict structure-property relationships between changes in chemical structure and changes in electronic structure. In an industrial application, the R2 correlation and the P-value are key metrics for evaluating the performance of computational chemistry methods. Across these materials, both iQCC and iQCC +PT results had the highest performance (Fig. 4, and Table 2), based on linear regression metrics R2, P-value, Pearson Correlation, and mean absolute error (MAE). The statistical analysis of the DFT based methods show mixed results, where the best performing DFT functionals based on MAE, tend to perform worse with respect to the linear fit parameters (R2 and P-value shown in Table 2). The T S0 estimates from iQCC alone are less accurate than those from iQCC +PT, underscoring the need for a PT correction unless the full set of entanglers is variationally optimized. The ansatz used for the PT correction contains the entire list of all entanglers that have a non-zero gradient (which allows for near arbitrary excitation operators that appear as qubit entanglers that can act upon all the qubits in the Hamiltonian) whereas the variationally bounded iQCC result is manually limited as to allow the optimization and normalization of the computed energies [19].
| Method | P-value | Pearson Corr. | MAE [eV] | Std. Dev. [eV] | |
| RHF/ROHF | 0.7393 | 8.04 | 0.8598 | 0.3809 | 0.1456 |
| RO-B3LYP | 0.5068 | 4.29 | 0.7119 | 0.1974 | 0.1604 |
| RO-CAM-B3LYP | 0.6085 | 9.98 | 0.7801 | 0.1161 | 0.1035 |
| RO-PBEh | 0.3062 | 4.01 | 0.5533 | 0.1457 | 0.1512 |
| RO-B97X | 0.8481 | 2.97 | 0.9209 | 0.2194 | 0.0843 |
| TD-B3LYP | 0.6638 | 3.86 | 0.8148 | 0.1209 | 0.0592 |
| TD-CAM-B3LYP | 0.8103 | 1.15 | 0.9001 | 0.2555 | 0.1029 |
| CCSD | 0.8070 | 1.28 | 0.8983 | 0.2200 | 0.1058 |
| CR-CC(2,3) | 0.7791 | 2.91 | 0.8827 | 0.2912 | 0.1130 |
| iQCC | 0.8786 | 7.61 | 0.9374 | 0.1180 | 0.0788 |
| iQCC+PT | 0.9411 | 9.69 | 0.9701 | 0.0501 | 0.0378 |
The iQCC and iQCC +PT are generally blue shifted from the experimental value, while CCSD and CR-CC(2,3) were both consistently red shifted, as shown in Fig. 4, which is in agreement with other observations [49]. This may indicate that a triplet state is more strongly correlated and CCSD and CR-CC(2,3) may “overcorrect” since they are not, in general, variationally bounded. The better MAE performance of the iQCC and iQCC +PT compared to the CCSD and CR-CC(2,3) likely results from the fact that the approximate QCC ansatz provides a more balanced description of a wave function by dropping unimportant triple- and disconnected quadruple-excitations while incorporating necessary higher excitation operators. The iQCC results are variationally bounded and while the entire QCC ansatz was not simulated the iQCC results are objectively better than CCSD and CR-CC(2,3) based on the statistics (Table 2). This demonstrates that despite CR-CC(2,3) being a common “gold standard” benchmark [45, 46, 47], it would not be a reliable benchmark for quantum algorithms in the field of phosphorescent emitters.
These results demonstrate that iQCC +PT not only achieves unprecedented classical emulation scale but also delivers predictive accuracy of phosphorescent T S0 gaps in line with, or surpassing, the best available classical methods. By reliably capturing structure–property trends in industrially relevant phosphorescent emitters, iQCC +PT establishes itself as a practical design tool for next-generation organo-metallic phosphorescent materials and a benchmark for evaluating quantum advantage in computational chemistry.
3 Discussion
The results here demonstrate that the iQCC algorithm if it were to be run on a fully error corrected universal quantum computer would provide a significant advantage in estimating the T S0 gap of Ir (III) and Pt (II) phosphorescent emitters when compared to established classical methods. The correlation plot between the experimental and simulated results for iQCC and iQCC +PT showed R2 of 0.87 and 0.94 respectively, surpassing the best performing conventional methods evaluated. This computational demonstration challenges the conjecture that because there are near linear scaling methods for coupled cluster that can simulate very large molecules, that using a quantum computer to solve for ground state energies would not provide an advantage over conventional methods.
The results here show promise for using the iQCC in a materials discovery process for new OLED phosphorescent emitters, specifically Ir (III) and Pt (II) emitters, which are not considered strongly correlated. In future work, we aim to study systems that are known to be strongly correlated and compare against DMRG benchmarks. Additionally, we note that the iQCC quantum-on-classical-hardware simulations are not limited to the simulation of the T S0 transition, and with the future improvements to the algorithm we intend to further push the quantum simulation boundary, by considering both strong correlation and multireference character. Ultimately, we aim to position iQCC as a competitive tool for quantum chemistry simulations, which provides both efficient and accurate results on a wide range of chemical systems.
Our solver acts as a production tool today, delivering quantum-algorithm results without quantum hardware or expensive supercomputing clusters. Tomorrow, it becomes a benchmarking tool, providing the gold-standard outputs against which quantum computers running chemistry/materials/drug-discovery problems can be validated.
4 Methods and Materials
4.1 Qubit Hamiltonian and QCC ansatz
The QCC method starts from the second-quantized electronic Hamiltonian of a molecule:
constructed from fermionic creation and annihilation operators and . The coefficients and are values of one- and two-electron integrals written in a spin-orbital basis.
The electronic Hamiltonian above is constructed using classical methods such as restricted Hartree–Fock (RHF) or restricted open-shell Hartree–Fock (ROHF) to generate the fermionic Hamiltonian, which is expressed in qubit form as
where each is a Pauli word (a tensor product of , , , or operators), and are real coefficients obtained by applying the Jordan–Wigner (JW) transformation to the usual second-quantized fermionic Hamiltonian [50]. This qubit Hamiltonian forms the basis for the qubit coupled cluster (QCC) method.
The QCC ansatz is constructed as a product of one-parameter unitary entangling operations,
where each is a Hermitian Pauli-word generator (an “entangler”), and each is an amplitude to be optimized. The QCC energy for a fixed qubit mean-field reference (in this paper, the Hatree-Fock reference state was used) state is
4.2 Iterative QCC and generator selection
The iQCC framework constructs the QCC ansatz by adding generators that most effectively lower the energy. For a candidate generator, the zero-amplitude energy gradient is
Generators with large provide the steepest local descent directions.
To avoid evaluating gradients for all possible Pauli words, iQCC employs the direct interaction set (DIS), defined as the subset of entanglers guaranteed to yield non-zero gradients. Membership in the DIS is determined by algebraic criteria based on the structure of the Hamiltonian and the action of .
After optimizing a new amplitude for a generator , the Hamiltonian is updated via the dressing transformation as,
This operation refines the Hamiltonian while preserving its spectrum. The dressing step increases the number of Hamiltonian terms; truncation of small coefficients controls the effective Hamiltonian size.
Outline of the iQCC iteration.
Each iQCC iteration consists of the following steps:
1.
Generate entanglers: Construct the DIS of Pauli-word generators satisfying the algebraic criteria for non-zero gradients. Evaluate (or estimate) the zero-amplitude gradients , rank all DIS generators by , and select the highest-impact ones to promote to entanglers.
2.
Optimize amplitudes: Refine the amplitudes associated with the newly added entanglers within a reduced variational subspace [26].
3.
Dress the Hamiltonian: Apply the unitary dressing transformation using the optimized amplitudes, and truncate Pauli terms whose coefficients fall below the numerical precision threshold.
4.
Evaluate convergence: Repeat the iteration until both the energy and amplitude changes satisfy the stopping criteria.
To target specific spin manifolds (e.g., the triplet state), we augment the variational objective with an penalty functional [51],
where is the JW-transformed total spin operator and is a penalty parameter (penalizing either the deviation from number of electrons or spin state). A full derivation and further details on the QCC ansatz and iQCC procedure from a theoretical perspective are outlined in Ref [18].
4.3 Parallel iQCC implementation
The novelty of the parallel iQCC approach lies in its partitioning strategy: Pauli words are expressed in binary representation, and selected bits are used as partitioning keys [27]. This provides a clear mapping of where any given Pauli word resides across CPUs. This mapping is particularly critical in the dressing step, where new Pauli words are generated. To maintain efficiency, newly generated terms must be checked against existing ones and combined when appropriate; otherwise, each CPU risks redundantly reconstructing the full Hamiltonian, negating the benefits of parallelization. With the partitioning scheme in place, we can avoid costly all-to-all communication—where every CPU must broadcast its new Pauli words to every other CPU. Since inter-CPU communication is itself a major computational resource, minimizing it is essential. The design of parallel iQCC must therefore satisfy two constraints: (i) distributing Pauli words across CPUs to prevent memory overload on any single node, and (ii) reducing CPU-to-CPU communication to the absolute minimum (saving the time for communication among cpus). Balancing these requirements is key to extending the tractability of iQCC simulations on classical HPC platforms. The partitioning scheme is based on selecting a bitmask that identifies which bits in the binary representation of a Pauli word are used for CPU assignment. The rank of a Pauli word under this mask is computed by packing the selected bits into an integer, which then determines the CPU to which the Pauli word is assigned. This method ensures that each CPU can independently determine the destination of newly generated Pauli words during the dressing step, using simple bitwise operations without requiring extensive communication between CPUs.
A major advantage of this scheme is that Pauli multiplication becomes simple bitwise XOR, allowing each processor to locally determine where newly generated terms belong without global searches or all-to-all communication. This dramatically reduces inter-CPU traffic—one of the dominant costs in large-scale simulations—and distributes memory evenly across nodes. As a result, the parallel C++ iQCC implementation achieves significant speedup in dressing and enables efficient optimization of both exact and approximate QCC ansätze at much larger system sizes than previously feasible.
Applying iQCC to large active spaces requires parallelization of both dressing and amplitude optimization. We implemented a distributed-memory parallel version of iQCC in C++ with OpenMPI, based on a bit-partitioning strategy for distributing Pauli words across processors.
Each Pauli word acting on qubits is encoded in a symplectic binary vector
where and encode the presence of and operators, respectively. Pauli multiplication corresponds to bitwise XOR on this representation.
A processor assignment is defined using a bitmask that selects positions of contributing to the rank,
where . A Pauli word is mapped to processor .
The essential algebraic property enabling efficient parallelization is
where denotes bitwise XOR. Thus, during dressing, each processor can locally determine the destination of any newly generated Pauli word—no all-to-all communication or global lookup tables are required. This eliminates communication bottlenecks, balances memory across nodes, and enables strong scaling of iQCC (see the Results Section 2.2).
4.4 Perturbative a posteriori energy corrections
At each iteration of the iQCC procedure, the dressed Hamiltonian defines a direct interaction set (DIS) of generators. Each DIS generator corresponds to a mean-field excited configuration connected to the qubit mean-field reference state . Two quantities required for perturbation theory are computed during the DIS construction:
1.
The gradient magnitude
(14) which corresponds to the first energy derivative associated with generator .
2.
The excited-state energy
(15) where the second equality uses the factorization to refer to unique basis states .
Defining the reference energy and the energy denominator
the iQCC implementation evaluates two classical a posteriori energy corrections.
The standard EN2 expression is
yielding the corrected energy
4.5 Simulation Protocols
The 16 qubit nitrogen molecule simulation was conducted at R=2.05 A with cc-pVDZ basis with singlet configuration [26]. The 36 qubit water molecule simulation was conducted using equilibrium geometry ((O-H) = 0.95 Å and HOH at 107.6°) and 6-31G(d) basis [52]. The CAS was generated by using all RHF molecular orbitals (MOs) except the lowest-energy MO, which is the core 1s orbital of the oxygen atom, resulting in CAS(8,18) [26]. Here, the ground state energy was computed using 2000 entanglers with 2nd order polynomial QCC ansatz.
The singlet and triplet geometries were optimized using either Gaussian [53] or TeraChem [54, 55] with LANL2DZ-ECP [56, 57, 58] on the metal centers and 6-31G** [41, 24, 28, 20, 26] on all other atoms, combined with the B3LYP DFT functional in the unrestricted formalism [4]. DFT energies were computed using two methods: TD-DFT from the singlet geometry with an S T1 transition employing the B3LYP and CAM-B3LYP functionals [59, 4, 60], and the T S0 energy difference in the triplet geometry (using the triplet optimized geometry for both T1 and S0 electron configuration) with restricted open-shell DFT using the B3LYP, CAM-B3LYP, PBEh, and B97X functionals [4, 60].
The active space used follows the same method described in [28] with variable CAS sizes. Two Ir complexes (Q2 and Q4) that exhibited poor performance at CAS(70,70) were increased to CAS(100,100) to test whether the CAS size was insufficient. Upon completion, it was determined that CAS(100,100) provided much better agreement with experiment compared to CAS(70,70), so the iQCC, CCSD, and CR-CC(2,3) used the larger CAS for statistical analysis. To double-check the iQCC results, CCSD and CR-CC(2,3) calculations were also performed using the same one- and two-electron integrals generated by GAMESS [61, 62].
Molecular Hamiltonians in qubit basis were prepared using the triplet geometry of the materials by running triplet ROHF and singlet RHF to generate the 1 and 2 electron integrals in molecular orbital basis using a modified version of GAMESS [61, 62], followed by transformation into qubit basis using the Jordan-Wigner transformation [63]. The selection criteria were to select a fixed number of occupied orbitals below the Fermi level and the same number of virtual orbitals above [28]. Two Ir complexes Q2 and Q4 which had poor performance at CAS(70,70) were increased to CAS(100,100) to check if the CAS was insufficient, and upon finishing the computation it was determined that CAS(100,100) had a much better agreement with experiment compared to CAS(70,70). To double check the iQCC results, CCSD and CR-CC(2,3) was also run using the same 1 & 2 electron integrals using GAMESS [43, 44]. CISD energies were only run for Q1 to determine the number of entanglers required for the final iQCC step. The CISD energies for the T1 state at CAS(100,100) could not be converged.
The iQCC algorithm was run using the similar approach to the previous literature [51, 28, 26]. In this instance, there were 8 iQCC dressing steps that contained 28 entanglers of which the entanglers were optimized to the 6th order, followed by 6 iQCC steps with 46 entanglers optimized at the 4th order. This was then followed by a repeating procedure that would optimize 300 entanglers at the 2nd order, but only dress 46 entanglers per iQCC step until the amplitudes of all 300 entanglers was less than 0.012 at which a final iQCC step was conducted using a set number of entanglers at the 1st order depending on the CAS size (300,000 entanglers for CAS(40,40), 400,000 entanglers for CAS(50,50), 600,000 entanglers for CAS(70,70), 800,000 entanglers for CAS(80,80), and 1,500,000 entanglers for CAS(100,100)). The iQCC simulations for the OLED phosphorescent emitters were run on compute nodes using two AMD EPYC 7702 64-core processors with 64 processes.
4.6 Synthesis and Measurement
Seven Ir (III) and five Pt (II) phosphorescent materials were either synthesized using published synthetic routes [23, 24, 38, 39, 40] or purchased. The PL spectra were measured using Hamamatsu Quantaurus-QY C11347 spectrometer at 77 K. Materials Q1, Q5, Q9, Q13, Q6, Q10, Q3, Q4, Q2 were measured in toluene, Q7 was measured in tetrahydrofuran, and Q8 and Q11 were measured in dichloromethane. The literature values for the peak PL emission of Q12 and Q14 were reported to be measured in 2-methyltetrahydrofuran at 77 K [23, 24].
5 Acknowledgments
The authors would like to acknowledge the funding provided by Innovation Solutions Canada project #202208-F0033-C00003 and Next Generation Manufacturing Grant #14234. The authors would like to thank Hyun-Ju Lim at the Embassy of Canada in the Republic of Korea for helping facilitate the collaboration through the Canadian International Innovation Program.